OpenStudy (anonymous):

solve x^2-10x=1575 by completing the square.

OpenStudy (anonymous):

x^2-10x=1575 add 25 to each sides x^2-10x+25=1600 (x-5)^2=1600 (x-5)^2=40^2 x=45 or x=-35

OpenStudy (anonymous):

first move 1575 to the left hand side x^2-10x-1575=0 next complete the square so we have (x-5)^2 -25 - 1575 = 0 (x-5)^2 - 1600 = 0